# Incremental Generation is Necessity and Sufficient for Universality in Flow-Based Modelling

### v1

@@ 3.1 Negative Results: Incremental Generation is Necessary
Our first main results shows that the collection of all flowable homeomorphisms compactly supported on $\delta,1-\delta]^{d}$ for any given $0<\delta<\tfrac{1}{2}$. (a subset of $\operatorname{\mathcal{H}}_{d}([\delta,1-\delta]^{d})$) are a small subset of homeomorphisms compactly supported on $0,1]^{d}$. This means there is not only one but many functions in $\operatorname{\mathcal{H}}_{d}([0,1]^{d})$ which cannot be approximated by flowable functions in $\operatorname{\mathcal{H}}_{d}([\delta,1-\delta]^{d})$ for any $0<\delta<\tfrac{1}{2}$.
@@
Figure 2: Why Non-Incremental Generation is Not Universal: The reason why non-incremental generators (Theorem 1) fail to be universal is that most homeomorphisms cannot be represented as flows (Theorem 2). The idea is that there is a dense open set of orientation-preserving homeomorphisms supported on the hypercube $0,1]^{d}$, which can be approximated/perturbed so that any given orbit becomes periodic (SubFigure 2(a)). Then, these perturbations can be further perturbed so that the a small neighbourhood around the given orbit becomes a basin of attraction (SubFigure 2(b)), which cannot happen for any flow. Consequently, the complement of any such map, which contains the set $\mathrm{NODE}_{\sigma}^{\star}([0,1]^{d})$ cannot be dense, implying that all non-incremental generators/autonomous Neural ODEs fail to be universal approximators of orientation-preserving homeomorphisms supported on the hypercube $0,1]^{d}$.
Theorem 1 (Non-Incremental Generation is Not Universal)
Let $d\in\mathbb{N}_{+}$ with $d>1$ and consider the class of all non-incremental generators, i.e. autonomous neural ODEs, for any Lipschitz activation function
[…]
Then, $\mathrm{NODE}_{\sigma}((0,1)^{d})$ is nowhere dense in $\operatorname{\mathcal{H}}_{d}([0,1]^{d})$.
Now, Theorem 1 is implied by our more general result in topological dynamics, illustrated in Figure 2,
@@
Let $d\in\mathbb{N}_{+}$ with $d>1$. The set of flowable homeomorphisms in $\operatorname{\mathcal{H}}_{d}([0,1]^{d})$ is meager in $\operatorname{\mathcal{H}}_{d}(P)$ for every $d>1$ and compact set $P$ s.t. $0,1]^{d}\subset int(P)$.
@@ 3.2.1 Qualitative Universal Approximation Guarantee
Let $d\in\mathbb{N}_{+}$ and $d\geq 5$. There exists a constant $K_{d}\in\mathbb{N}_{+}$ such that: for every $\varphi\in\operatorname{\mathcal{H}}_{d}([0,1]^{d})$ and every $\varepsilon>0$ there exists some $G\leq K_{d}$ and ReLU neural ODEs $\{\Psi^{(g)}=\mathrm{Flow}(\Phi^{(g)})\}_{g=1}^{G}\subset\mathcal{X}^{0}$ such that the diffeomorphism $\Psi\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\bigcirc_{g=1}^{G}\,\Psi^{(g)}$ is Lipschitz and compactly supported and satisfies the approximation guarantee:
@@
Figure 3: By (a uniform version) of Thursten’s Theorem thurston1974foliations we $\operatorname{\mathcal{H}}_{d}([0,1]^{d})$ is a simple group and since the group generated by flows is normal, then every diffeomorphism (green curve) must be the composition of finitely many flows of vector fields $V^{(1)}$, $\dots$, $V^{(G)})$ (here $G=4$) – Sub-Figure 3(a). Each vector field is then approximated by a ReLU MLP at an optimal rate (Sub-Figure 3(b)) with maximal Lipschitz regularity using Hong_2024__ReLUMLPs; the approximation of the original homeomorphism (Theorem 7) is concluded using Grönwall’s inequality.
@@ 3.2.2 Quantitative Formulations
We start this section by remarking that, the class $\operatorname{\mathcal{H}}_{d}^{G,0}([0,1]^{d})$ for some $G>1$ is non-empty. Thus the theorem provides a non-vacuous statement for $G>1$.
@@
there exists some $G\in\mathbb{N}_{+}$ and $G>1$ that $\operatorname{\mathcal{H}}_{d}^{G,0}([0,1]^{d})$ is non-empty,
@@
Let $n,d\in\mathbb{N}$, and $\varphi\in\operatorname{\mathcal{H}}_{d}^{G,0}([0,1]^{d})$ then, there exists ReLU neural ODEs $\{\Psi^{(g)}=\mathrm{Flow}(\Phi^{(g)})\}_{g=1}^{G}\subset\mathcal{X}^{0}$ such that $\Psi\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\bigcirc_{g=1}^{G}\,\Psi^{(g)}$ satisfies the approximation guarantee:
$\big\|\varphi-\Psi\big\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}\leq\sum_{g=1}^{G}\left(2\|\omega^{(g)}(\frac{d}{2n})\|_{l^{\infty}}\prod_{j=g}^{G}e^{L^{V^{(j)}}}\right)$ (9)
@@
Moreover, $\Psi$ is a compactly-supported homeomorphism on $\mathbb{R}^{d}$ supported on $(0,1)^{d}$, with Lipschitz constant at-most $\prod_{g=1}^{G}e^{L^{\Phi^{(g)}}}$ and $\Phi^{(1)},\dots,\Phi^{(G)}$ are ReLU MLP vector fields of depth $\lceil\text{log}_{2}d\rceil+6$, width $8d(n+1)^{d}+9$, and at-most $16d(n+1)^{d}+9$ non-zero parameters. In particular, these vector fields and $L^{\Psi}$ do not depend on the parameter $n$.
Furthermore, $\omega^{(g)}$ is the modulus of regularity of $\varphi^{(g)}$ which is equal to the modulus of regularity of $\Phi^{(g)}$.
@@
Let $\varphi\in\operatorname{\mathcal{H}}_{d}^{G,s}([0,1]^{d})$ and $N,\,L,\,s\in\mathbb{N}_{+}$ then, there exists ReLU neural ODEs $\{\Psi^{(g)}=\mathrm{Flow}(\Phi^{(g)})\}_{g=1}^{G}\subset\mathcal{X}^{0}$ such that $\Psi\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\bigcirc_{g=1}^{G}\,\Psi^{(g)}$ satisfies the approximation guarantee:
$\begin{split}\big\|\varphi-\Psi\big\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}\leq\sum_{g=1}^{G}\left(2\|\omega^{(g)}(N,L)\|_{l^{\infty}}\prod_{j=g}^{G}e^{L^{V^{(j)}}}\right)\end{split}$ (10)
@@
Moreover, $\Psi$ is a compactly-supported $C^{s}$-diffeomorphism on $\mathbb{R}^{d}$ supported on $(0,1)^{d}$, with Lipschitz constant at-most $\prod_{g=1}^{G}e^{L^{\Psi^{(g)}}}$ and $\Phi^{(1)},\dots,\Phi^{(G)}$ are ReLU MLPs of width less than or equal to $17s^{d+1}3^{d}d^{2}(N+2)\log_{2}(8N)$ and depth $18s^{2}(L+2)\log_{2}(4L)+2(d+1)$ where $N,L\in\mathbb{N}_{+}$. In particular, $L^{\Psi}$ depends on the parameter $N$.
@@
There exists a positive number $K_{d}\in\mathbb{N}_{+}$ such that any diffeomorphism $\varphi\in\operatorname{Diff}_{0}([0,1]^{d})$ can be written as at most $K_{d}$ flows.
@@
Let $\varphi\in\operatorname{Diff}_{0}([0,1]^{d})$ and $N,L\in N_{+}$ and take $L$ fixed; then, for $\forall s\in\mathbb{N}_{+}$ there exists ReLU neural ODEs $\{\Psi^{(g)}=\mathrm{Flow}(\Phi^{(g)})\}_{g=1}^{G}\subset\mathcal{X}^{0}$ such that $\Psi\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\bigcirc_{g=1}^{G}\,\Psi^{(g)}$ satisfies the approximation guarantee:
@@
Moreover, $\Psi$ is a compactly-supported $C^{s}$-diffeomorphism on $\mathbb{R}^{d}$ supported on $(0,1)^{d}$, with Lipschitz constant at-most $\prod_{g=1}^{G}e^{L^{\Psi^{(g)}}}$ and $\Phi^{(1)},\dots,\Phi^{(G)}$ are ReLU MLPs of width less than or equal to $17s^{d+1}3^{d}d^{2}(N+2)\log_{2}(8N)$ and depth $18s^{2}(L+2)\log_{2}(4L)+2(d+1)$ where $N,L\in\mathbb{N}^{+}$. In particular, $L^{\Psi}$ depends on the parameter $n$.
@@ 4.1 Universal Approximation of arbitrary Lipschitz functions between arbitrary dimensions
$\frac{dx_{t}^{(x,y)}}{dt}=0,\qquad\frac{dy_{t}^{(x,y)}}{dt}=f(x_{t}^{(x,y)}),\mbox{ where }(x_{0},y_{0})=(x,0)$
@@
$\frac{dx_{t}^{(x,y)}}{dt}=0\Rightarrow x_{t}^{(x,y)}=x\Rightarrow y_{t}^{(x,y)}=y_{0}+\int_{0}^{t}f(x)dt\Rightarrow y_{t}^{(x,y)}=0+t\,f(x)$
@@
Let $d,D\in\mathbb{N}_{+}$, $L\geq 0$, and $f:\mathbb{R}^{d}\to\mathbb{R}^{D}$ be Lipschitz. For every $n\in\mathbb{N}_{+}$ and each $i\in[D]$ there exists ReLU Neural ODEs $\Psi_{i}:\mathbb{R}^{d+1}\to\mathbb{R}^{d+1}$ such that the Latent Neural ODE $\Psi\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\bigoplus_{i=1}^{D}(\pi_{1}^{d+1}\circ\Psi_{i}\circ\iota_{d}^{d+1})$ satisfies
$\big\|f(x)-\Psi\big\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}\in O\big(\tfrac{1}{n}\big).$
@@
Thus, Corollary 11 provides quantitative and, more topologically more explicit, version of the very recent qualitative result of de2025approximation as a direct consequence of our main theorem.
@@
The question of the optimality of the rate $\mathcal{O}(n^{d+1})$ is not known, as there are no available approximation theoretic lower bounds, nor tools for establishing lower bounds, in our non-vector space setting it is currently unknown. However, is likely that a rate of $\mathcal{O}(n^{d})$ cannot be achieved while requiring that the core of the model is the conjugation of a homeomophism by linear maps, due to the need to lift in order to approximate general continuous functions.
@@
We examine the implications our results to classical generative modelling, in the sense of Wasserstein GANs arjovsky2017wasserstein; korotin2021wasserstein, and highlighting the differences and similarities.
@@ 4.2 Universal Approximation of measures satisfying Caffarelli Conditions
Let $\mu$ and $\nu$ be measures which are absolutely continuous with respect to the uniform measure $U$ on $(0,1)^{d}$; and whose Radon-Nikodym densities $\tfrac{d\mu}{dU}$ and $\tfrac{d\nu}{dU}$ are $\alpha$-Hölder for some $\alpha\in(0,1)$, and are bounded (above and below on $0,1]^{d}$).
@@
Let $2<d\in\mathbb{N}_{+}$ and $\mu,\nu$ be probability measures on $0,1]^{d}$ satisfying Assumption 4.1. For every $\varepsilon,\delta>0$, there exists a large enough $N\in\mathbb{N}_{+}$ and a Latent Neural ODE $\Psi:\mathbb{R}^{d+D}\to\mathbb{R}^{d+D}$ such that for all i.i.d. random variables $Z_{1}\sim\dots\sim Z_{N}$, with law $\nu$ (defined on a common probability space $(\Omega,\mathcal{F},\mathbb{P})$)

### v2

@@ 3.1 Negative Results: Incremental Generation is Necessary
Our first main result shows that the collection of all flowable homeomorphisms compactly supported on $0,1]^{d}$ is non-universal in $\operatorname{\mathcal{H}}_{d}([0,1]^{d})$; moreover, the collection of all flowable homeomorphisms compactly supported on $\delta,1-\delta]^{d}$ for any given $0<\delta<\tfrac{1}{2}$ are a small subset of homeomorphisms compactly supported on $0,1]^{d}$. This means there is not only one but many functions in $\operatorname{\mathcal{H}}_{d}([0,1]^{d})$ which cannot be approximated by flowable functions in $\operatorname{\mathcal{H}}_{d}([\delta,1-\delta]^{d})$ for any $0<\delta<\tfrac{1}{2}$.
@@
Figure 2: Why Non-Incremental Generation is Not Universal: The reason why non-incremental generators (Theorem 2) fail to be universal is that most homeomorphisms cannot be represented as flows (Theorem 3). The idea is that there is a dense open set of orientation-preserving homeomorphisms supported on the hypercube $0,1]^{d}$, which can be approximated/perturbed so that any given orbit becomes periodic (SubFigure 2(a)). Then, these perturbations can be further perturbed so that the a small neighbourhood around the given orbit becomes a basin of attraction (SubFigure 2(b)), which cannot happen for any flow. Consequently, the complement of any such map, which contains the set $\mathrm{NODE}_{\sigma}((0,1)^{d})$ cannot be dense, implying that all non-incremental generators/autonomous Neural ODEs fail to be universal approximators of orientation-preserving homeomorphisms supported on the hypercube $0,1]^{d}$.
Theorem 1 (Incremental Generation is Necessary for Universality)
$\mathrm{NODE}_{\sigma}([0,1]^{d})$ is not universal in $\operatorname{\mathcal{H}}_{d}([0,1]^{d})$.
Theorem 2 (Non-Incremental Generation on $\delta,1-\delta]^{d}$ is meagre)
Let $d\in\mathbb{N}_{+}$ with $d>1$ then for any Lipschitz activation function $\mathrm{NODE}_{\sigma}([\delta,1-\delta]^{d})$ is nowhere dense in $\operatorname{\mathcal{H}}_{d}([0,1]^{d})$.
Theorem 2 is implied by our more general result in topological dynamics, illustrated in Figure 2,
@@
Let $d\in\mathbb{N}_{+}$ with $d>1$. The set of flowable homeomorphisms in $\operatorname{\mathcal{H}}_{d}([0,1]^{d})$ is meagre in $\operatorname{\mathcal{H}}_{d}(P)$ for every $d>1$ and compact set $P$ s.t. $0,1]^{d}\subset int(P)$.
@@ 3.2.1 Qualitative Universal Approximation Guarantee
Let $d\in\mathbb{N}_{+}$ and $d\geq 5$. There exists a constant $T_{d}\in\mathbb{N}_{+}$ such that: for every $\varphi\in\operatorname{\mathcal{H}}_{d}([0,1]^{d})$ and every $\varepsilon>0$ there exists some $T\leq T_{d}$ and ReLU neural ODEs $\{\Psi^{(t)}=\mathrm{Flow}(\Phi^{(t)})\}_{t=1}^{T}\subset\mathcal{X}^{0}$ such that the homeomorphism $\Psi\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\bigcirc_{t=1}^{T}\,\Psi^{(t)}$ is Lipschitz and compactly supported and satisfies the approximation guarantee:
@@
Figure 3: By (a uniform version) of Thursten’s Theorem thurston1974foliations we $\operatorname{\mathcal{H}}_{d}([0,1]^{d})$ is a simple group and since the group generated by flows is normal, then every diffeomorphism (green curve) must be the composition of finitely many flows of vector fields $V^{(1)}$, $\dots$, $V^{(T)})$ (here $T=4$) – Sub-Figure 3(a). Each vector field is then approximated by a ReLU MLP at an optimal rate (Sub-Figure 3(b)) with maximal Lipschitz regularity using Hong_2024__ReLUMLPs; the approximation of the original homeomorphism (Theorem 8) is concluded using Grönwall’s inequality.
@@ 3.2.2 Quantitative Formulations
We start this section by remarking that, the class $\operatorname{\mathcal{H}}_{d}^{T,0}([0,1]^{d})$ for some $T>1$ is non-empty. Thus the theorem provides a non-vacuous statement for $T>1$.
@@
there exists some $T\in\mathbb{N}_{+}$ and $T>1$ such that $\operatorname{\mathcal{H}}_{d}^{T,0}([0,1]^{d})$ is non-empty,
@@
Let $n,d\in\mathbb{N}$, and $\varphi\in\operatorname{\mathcal{H}}_{d}^{T,0}([0,1]^{d})$ then, there exists ReLU neural ODEs $\{\Psi^{(t)}=\mathrm{Flow}(\Phi^{(t)})\}_{t=1}^{T}\subset\mathcal{X}^{0}$ such that $\Psi\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\bigcirc_{t=1}^{T}\,\Psi^{(t)}$ satisfies the approximation guarantee:
$\big\|\varphi-\Psi\big\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}\leq\sum_{t=1}^{T}\left(2\|\omega^{(t)}(\frac{d}{2n})\|_{l^{\infty}}\prod_{j=t}^{T}e^{L^{V^{(j)}}}\right)$ (9)
@@
Moreover, $\Psi$ is a compactly-supported homeomorphism on $\mathbb{R}^{d}$ supported on $(0,1)^{d}$, with Lipschitz constant at-most $\prod_{t=1}^{T}e^{L^{\Phi^{(t)}}}$ and $\Phi^{(1)},\dots,\Phi^{(T)}$ are ReLU MLP vector fields of depth $\lceil\text{log}_{2}d\rceil+6$, width $8d(n+1)^{d}+9$, and at-most $16d(n+1)^{d}+9$ non-zero parameters. In particular, these vector fields and $L^{\Psi}$ do not depend on the parameter $n$.
Furthermore, $\omega^{(t)}$ is the modulus of regularity of $\varphi^{(t)}$ which is equal to the modulus of regularity of $\Phi^{(t)}$.
Neural ODE $\Psi$ is compactly supported on $(0,1)^{d}$ showing that incrementality breaks meagreness of $\mathrm{NODE}([\delta,1-\delta]^{d})$ for any $\delta\in(0,\frac{1}{2})$.
@@
Let $\varphi\in\operatorname{\mathcal{H}}_{d}^{T,s}([0,1]^{d})$ and $N,\,L,\,s\in\mathbb{N}_{+}$ then, there exists ReLU neural ODEs $\{\Psi^{(t)}=\mathrm{Flow}(\Phi^{(t)})\}_{t=1}^{T}\subset\mathcal{X}^{0}$ such that $\Psi\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\bigcirc_{t=1}^{T}\,\Psi^{(t)}$ satisfies the approximation guarantee:
$\begin{split}\big\|\varphi-\Psi\big\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}\leq\sum_{t=1}^{T}\left(2\|\omega^{(t)}(N,L)\|_{l^{\infty}}\prod_{j=t}^{T}e^{L^{V^{(j)}}}\right)\end{split}$ (10)
@@
Moreover, $\Psi$ is a compactly-supported homeomorphism on $\mathbb{R}^{d}$ supported on $(0,1)^{d}$, with Lipschitz constant at-most $\prod_{t=1}^{T}e^{L^{\Psi^{(t)}}}$ and $\Phi^{(1)},\dots,\Phi^{(T)}$ are ReLU MLPs of width less than or equal to $17s^{d+1}3^{d}d^{2}(N+2)\log_{2}(8N)$ and depth $18s^{2}(L+2)\log_{2}(4L)+2(d+1)$ where $N,L\in\mathbb{N}_{+}$. In particular, $L^{\Psi}$ depends on the parameter $N$.
@@
There exists a positive number $T_{d}\in\mathbb{N}_{+}$ such that any diffeomorphism $\varphi\in\text{Diff}_{0}([0,1]^{d})$ can be written as at most $T_{d}$ flows.
@@
Let $\varphi\in\operatorname{Diff}_{0}([0,1]^{d})$ and $N,L\in N_{+}$ and take $L$ fixed; then, for $\forall s\in\mathbb{N}_{+}$ there exists ReLU neural ODEs $\{\Psi^{(t)}=\mathrm{Flow}(\Phi^{(t)})\}_{t=1}^{T}\subset\mathcal{X}^{0}$ such that $\Psi\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\bigcirc_{t=1}^{T}\,\Psi^{(t)}$ satisfies the approximation guarantee:
@@
Moreover, $\Psi$ is a compactly-supported homeomorphism on $\mathbb{R}^{d}$ supported on $(0,1)^{d}$, with Lipschitz constant at-most $\prod_{t=1}^{T}e^{L^{\Psi^{(t)}}}$ and $\Phi^{(1)},\dots,\Phi^{(T)}$ are ReLU MLPs of width less than or equal to $17s^{d+1}3^{d}d^{2}(N+2)\log_{2}(8N)$ and depth $18s^{2}(L+2)\log_{2}(4L)+2(d+1)$ where $N,L\in\mathbb{N}^{+}$. In particular, $L^{\Psi}$ depends on the parameter $n$.
@@ 4.1 Universal Approximation of arbitrary Lipschitz functions between arbitrary dimensions
$\frac{dx_{t}^{(x,y)}}{dt}=0,\qquad\frac{dy_{t}^{(x,y)}}{dt}=f(x_{t}^{(x,y)}),\mbox{ where }(x_{0},y_{0})=(x,y)$
@@
$\frac{dx_{t}^{(x,y)}}{dt}=0\Rightarrow x_{t}^{(x,y)}=x\Rightarrow y_{t}^{(x,y)}=y_{0}+\int_{0}^{t}f(x)dt\Rightarrow y_{t}^{(x,y)}=y+t\,f(x)$
@@
Let $d,D\in\mathbb{N}_{+}$, $L\geq 0$, and $f:\mathbb{R}^{d}\to\mathbb{R}^{D}$ be L-Lipschitz. For every $n\in\mathbb{N}_{+}$ and each $i\in[D]$ there exists ReLU Neural ODEs $\Psi_{i}:\mathbb{R}^{d+1}\to\mathbb{R}^{d+1}$ such that the Latent Neural ODE $\Psi\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\bigoplus_{i=1}^{D}(\pi_{1}^{d+1}\circ\Psi_{i}\circ\iota_{d}^{d+1})$ satisfies
$\big\|f(x)-\Psi\big\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}\in\mathcal{O}\big(\tfrac{1}{n}\big).$
@@
Thus, Corollary 12 provides a quantitative and, more topologically more explicit, version of the very recent qualitative result of de2025approximation as a direct consequence of our main theorem.
@@
The question of the optimality of the rate $\mathcal{O}(n^{d+1})$ is not known, as there are no available approximation theoretic lower bounds, nor tools for establishing lower bounds, in our non-vector space setting it is currently unknown. However, is likely that a rate of $\mathcal{O}(n^{d})$ cannot be achieved while requiring that the core of the model is the conjugation of a homeomorphism by linear maps, due to the need to lift in order to approximate general continuous functions.
@@
We examine the implications of our results to classical generative modelling, in the sense of Wasserstein GANs arjovsky2017wasserstein; korotin2021wasserstein, and highlight the differences and similarities.
@@ 4.2 Universal Approximation of measures satisfying Caffarelli Conditions
Let $\mu$ and $\nu$ be measures which are supported on $(0,1)^{d}$ and absolutely continuous with respect to the uniform measure $U$ on $(0,1)^{d}$; whose Radon-Nikodym densities $\tfrac{d\mu}{dU}$ and $\tfrac{d\nu}{dU}$ are $\alpha$-Hölder for some $\alpha\in(0,1)$, and are bounded (above and below on $(0,1)^{d}$).
@@
Let $2<d\in\mathbb{N}_{+}$ and $\mu,\nu$ be probability measures on $0,1]^{d}$ satisfying Assumption 4.1. For every $\varepsilon,\delta>0$, there exists a large enough $N\in\mathbb{N}_{+}$ and a Neural ODE $\Psi:\mathbb{R}^{d+D}\to\mathbb{R}^{d+D}$ such that for all i.i.d. random variables $Z_{1}\sim\dots\sim Z_{N}$, with law $\nu$ (defined on a common probability space $(\Omega,\mathcal{F},\mathbb{P})$)